# A physics-inspired momentum-based gradient method

### v1

@@ 3.2 Inverse Design of Multilayer Nanospheres
where N is the number of layers. Physical constraints include manufacturability (minimum layer thickness ¿5) and an upper bound on the overall radius $R_{\max}=300nm$
@@
where the lower bound ensures the validity of the effective continuous dielectric description and the upper bound reflects fabrication limits. The overall particle radius must not exceed a prescribed maximum value. For silver layers, the minimum thickness must be sufficiently large to form a continuous film. For $N=6$ layers, the optimized configuration yields an average absorptivity $J(x_{*})\approx 0.72$, compared to $\sim 0.45$ for a homogeneous $SiO_{2}$ sphere of the same radius. The optimal design exhibits a balance between plasmonic resonance enhancement from $Ag$ shells and dielectric confinement from $SiO_{2}$ layers. Figures 2 and 3 present the results. We can see that the proposed method converges much faster than the projected gradient descent, which becomes trapped in local minima.
Figure 2: Performance profiles of nonlinear momentum and projected gradient decent on the numerical example 2.
[…]
@@ 4 Conclusion
Based on the dynamics of non-Newtonian mechanical systems, this work proposes a novel nonlinear momentum optimization method. The method enhances the convergence speed and robustness of momentum-based gradient optimization algorithms by generalizing the kinetic energy form and introducing a nonlinear damping mechanism. Specifically, a anharmonic kinetic energy function is employed to characterize the inertial effect of accumulated gradient information during iterations, while the nonlinear damping adaptively regulates the influence of the momentum term on the convergence trajectory. Numerical experiments illustrate that, compared with classical momentum-based methods, the proposed algorithm exhibits notable advantages in parameter flexibility, stability, and robustness against gradient noise. Moreover, the method establishes new a physics-based perspective complementary to traditional numerical optimization analysis.
This study not only shows superior performance over conventional gradient methods in inverse photonic design tasks but also represents a new attempt to develop physics-inspired optimization algorithms. Future research directions include employing higher-order numerical integration schemes, introducing more general forms of kinetic energy, and combining advanced sampling and optimization strategies within Hamiltonian Monte Carlo frameworks to further improve efficiency. Additionally, systematic theoretical analysis is needed to deepen the understanding of the convergence and stability properties of nonlinear momentum mechanisms.

### v2

@@ 4.2 Inverse Design of Multilayer Nanospheres
where N is the number of layers. Physical constraints include manufacturability (minimum layer thickness ¿5nm) and an upper bound on the overall radius $R_{\max}=300nm$
@@
where the lower bound ensures the validity of the effective continuous dielectric description and the upper bound reflects fabrication limits. The overall particle radius must not exceed a prescribed maximum value. For silver layers, the minimum thickness must be sufficiently large to form a continuous film. For $N=3$ layers, the optimal design exhibits a balance between plasmonic resonance enhancement from $Ag$ shells and dielectric confinement from $SiO_{2}$ layers. Figure 3 and Figure 4 present the results. We can see that the Algorithm 1 and Algorithm 2 converge faster than HB, and AGD. Figure 4 also indicates that there are many local optima for the average absorption cross-section objective, which are related to different plasmonic resonance absorption spectrum structures for the multilayer nanosphere.
Figure 3: HB, AGD, and Algorithm 1 and Algorithm 2 for the numerical example 2. The learning rate $h$ 0.001, $\gamma h=0.1$ for all these algorithms. And $\eta=s=1.95$ for Algorithm 1 and Algorithm 2
Figure 4: Average absorption cross-sections over visible wavelength range. The method for computing the absorption cross section is mRTMA [20].
@@ 5 Conclusion
Based on non-Newtonian dynamics, this paper proposes a novel nonlinear momentum optimization method. The method significantly improves the convergence speed of momentum-based gradient optimization algorithms by generalizing the kinetic energy formulation and introducing a nonlinear damping mechanism. Specifically, an anharmonic kinetic energy function is used to capture the inertial effects of accumulated gradient information during iterations, while the nonlinear damping mechanism adaptively regulates the influence of the momentum term on the convergence trajectory. Numerical experiments show that the proposed algorithm offers significant advantages in terms of parameter flexibility, stability, and convergence speed compared to classical momentum-based methods. In particular, the method outperforms traditional gradient methods in inverse design tasks. Additionally, this method provides a new physics-based perspective on numerical optimization analysis, further expanding the application of physics-inspired optimization algorithms.
This study not only demonstrates significant performance improvements in inverse design tasks but also opens new paths for the development of physics-inspired optimization algorithms. Future research directions include: employing higher-order numerical integration schemes, exploring more general forms of kinetic energy, and combining advanced sampling and optimization strategies within the Hamiltonian Monte Carlo framework to further enhance efficiency. Additionally, systematic theoretical analysis is needed to deepen the understanding of the convergence and stability properties of nonlinear momentum mechanisms.